A munber is drawn at random from at hat that contain all the numbers from 1 to 100 .What is the prbability that the number is less than 16?
Answer by ewatrrr(24785)


Hi
100 numbers, of which 15 of them are < 16
P(number <16) = 15/100 = 3/20